OpenStudy (anikate):

What are rational and irrational numbers!!! HELP!! I am willing to stay up as long as you want me to. PLEASE HELP ME!! I'M IN A CRISIS!

OpenStudy (anikate):

how would 2/3 be a rational number? first of all, to get us started.

OpenStudy (chrisplusian):

rational numbers are ones that terminate or have a repeating strand. such as.25 (terminates) Or 2/3=.6666666666....... the 6 repeats so the strand has a repeating number. if the number does not terminate or have a repeating number then it is"irrational"

OpenStudy (anikate):

so 2/3 is rational right? how?

OpenStudy (chrisplusian):

yes

OpenStudy (anikate):

wha if it's\[-\sqrt{119} or \sqrt{-119 }\]

OpenStudy (anikate):

if 2/3 is rational why isn't pi irrational?

OpenStudy (chrisplusian):

if you have the square root of a negative # like the second example you gave you will have what is called an imaginary #

OpenStudy (anikate):

ok so what about the first one.

OpenStudy (anonymous):

rational number can be written as fraction. irrational number cannot be written as fraction, and usuallytend to be endless.

OpenStudy (anikate):

so how come .6 repeating is rational?

OpenStudy (anonymous):

.6 repeating can be written as 2/3, a fraction.

OpenStudy (chrisplusian):

It has a repeating sequence therefore it is rational

OpenStudy (anikate):

oh ok so can you plz give me examples of irrational numbers

OpenStudy (chrisplusian):

\[\sqrt{3}\]

OpenStudy (anikate):

ok

OpenStudy (anonymous):

irrational numbers such as pi, or square root of 2

OpenStudy (anikate):

is there an explaination that I can give for the sq rt. of 2 if it's rational or irrarional?

OpenStudy (anonymous):

square root of 2 cannot be written as fraction...

OpenStudy (anikate):

so it's irrational and Iwould write that my explaination is because the sq rt. of 2 cannot be written as a fraction?

OpenStudy (anonymous):

well, basically, yes.

OpenStudy (chrisplusian):

to fully understand it you have to think about what the deffinition of a square root is any number times itself has to equal the number under the square root symbol. you can get closer and closer to the two numbers that when multiplied by themself equl two but for every number you find that gets close there is yet one that is closer. so the possibilities are endless. and there for there is no logical sequence that when multiplied by itself will give you two

OpenStudy (anikate):

ok,so is there any other explanation?

OpenStudy (anikate):

虽然有i dont 个体他蛤他?

OpenStudy (anikate):

Sry i dont get that

OpenStudy (anikate):

but I get the point now

OpenStudy (anonymous):

then just say it's not a rational number because it cannot be written as fraction...

OpenStudy (anikate):

ok thx very much and everyone that helped plz just sau samething so I can give you a medal. :D :)